OpenStudy (anonymous):

Perform the requested operation or operations. f(x) = x-5/8 ; g(x) = 8x + 5, find g(f(x)).

OpenStudy (anonymous):

g(f(x))=8(x-5/8) + 5 = 8x

OpenStudy (anonymous):

i dont think thats it

OpenStudy (anonymous):

"g(f(x))" says compose g(x) of f(x), which means (literally) plug f(x) into g(x) for x in the equation g(x)

OpenStudy (anonymous):

well these are my anserws g(f(x)) = x - 5/8 g(f(x)) = x g(f(x)) = 8x + 35 g(f(x)) = x + 10

OpenStudy (anonymous):

is the problem f(x) = (x-5)/8, or f(x) = x- (5/8) ?

OpenStudy (anonymous):

(x-5)/8

OpenStudy (anonymous):

oh! well then my first response was wrong, sorry; g(f(x)) = 8((x-5)/8)+5 = ((8x-40)/8)+5 = (x-5)+5 = x

OpenStudy (anonymous):

thanks girlyyyy <3